# RMIT–ADM$+$S at the SIGIR 2025 LiveRAG Challenge

### v1

@@ Abstract.
This paper presents the RMIT–ADM+S participation in the SIGIR 2025 LiveRAG Challenge. Our Generation-Retrieval-Augmented Generation (GRAG) approach relies on generating a hypothetical answer that is used in the retrieval phase, alongside the original question. GRAG also incorporates a pointwise large language model (LLM)-based re-ranking step prior to final answer generation. We describe the system architecture and the rationale behind our design choices. In particular, a systematic evaluation using the Grid of Points (GoP) framework and $N$-way ANOVA enabled comparison across multiple configurations, including query variant generation, question decomposition, rank fusion strategies, and prompting techniques for answer generation. Our system achieved a Relevance score of 1.199 and a Faithfulness score of 0.477 on the private leaderboard, placing among the top four finalists in the LiveRAG 2025 Challenge.
@@
journalyear: 2025copyright: ccconference: ; July 17, 2025; Padua, Italybooktitle: SIGIR LiveRAG Challenge 2025, July 17, 2025, Padua, Italy
@@ 1. Introduction
Evaluation campaigns such as the SIGIR 2025 LiveRAG Challenge provide a structured and standardized setting for researchers and practitioners to develop and evaluate different Retrieval-Augmented Generation (RAG) approaches on a common dataset using shared metrics. These campaigns enable fair comparisons between systems in a controlled environment. LiveRAG 2025, the first edition of the challenge, required participants to develop RAG systems using the Falcon3-10B-Instruct model (Team, 2024) for final answer generation, thereby standardizing the generation component across all submissions. This constraint allows for a more focused evaluation of the retrieval components and prompts, as the generation model is fixed and does not introduce variability in the results. [^1] Other open-weight LLMs, such as the Llama models, were allowed in other components of the system, but only up to their 10B versions. In this context, the RMIT–ADM+S team submitted GRAG (Generation-Retrieval-Augmented Generation), a system selected through an internal evaluation process using a Grid of Points (GoP) approach and $N$-way ANOVA. [^2] The team includes members from the RMIT Centre for Human-AI Interaction (CHAI) and the ARC Centre of Excellence for Automated Decision-Making and Society (ADM+S).
Our submission builds on previous work and integrates several components: (1) hypothetical answer generation prior to retrieval; (2) large language model (LLM)-based query variant generation (Alaofi et al., 2023; Ran et al., 2025); (3) LLM-based re-ranking (Sun et al., 2025); and (4) answerability estimation in RAG systems (Pathiyan Cherumanal et al., 2024). We also developed supporting infrastructure for dynamic cloud-based resource allocation and LLM deployment in the AWS environment, enabling more efficient resource usage.
@@
The source code for both the system and the evaluation framework are publicly available at https://github.com/rmit-ir/GRAG-LiveRAG.
2. GRAG System Architecture
Figure 1 illustrates the pipeline of our proposed approach and the system used in our submitted run. In addition to the retrieval and answer generation stages of a RAG system, GRAG includes two additional stages: question augmentation and re-ranking. For all components involving LLMs, we used the same fixed open-weight model: Falcon3-10B-Instruct. The prompts used throughout the system are provided in Appendix A.
Figure 1. Pipeline of our GRAG approach. Components used in the final selected run are shown with solid borders; components analyzed during in-house evaluation are shown with dashed borders.
@@ 3.1.1. Dataset Preparation
Table 1. Configuration used to create the test collection using DataMorgana.
@@ 3.2. Grid of Points
Table 2. $N$-way ANOVA results. This analysis utilized the LLM-based evaluation approach, specifically the average-level evaluation method described in Section 3.1.3. A factor is considered to have a statistically significant impact if its probability value, PR(>F), is less than 0.05. The factors are presented in descending order based on their corresponding F-statistic values. The header lines are as follows: Factor: The factor being analyzed, refer to Section 3.1.2. SM: Sum of Squares. DF: Degrees of Freedom. F: F-statistic. PR(>F): Probability value greater than F, i.e., $p$-value.
Factor SM DF F PR(>F)
Question Augmentation* $0.3193$ $2$ $125.6920$ $<0.0001$
Query Variants Generation Prompts : Retrieval* $0.0062$ $1$ $4.9099$ $0.0297$
Query Variants Generation Prompts $0.0030$ $1$ $2.3892$ $0.1263$
Context Words Length Limits $0.0025$ $1$ $2.0052$ $0.1608$
Retrieval $0.0024$ $1$ $1.8631$ $0.1762$
Number of Documents Retrieved : Retrieval $0.0020$ $1$ $1.6100$ $0.2083$
Answer Generation Prompts $0.0014$ $1$ $1.1061$ $0.2962$
Context Words Length Limits : Question Augmentation $0.0008$ $2$ $0.3261$ $0.7227$
Context Words Length Limits : Answer Generation Prompts : Question Augmentation $0.0004$ $2$ $0.1566$ $0.8553$
Answer Generation Prompts : Question Augmentation $0.0004$ $2$ $0.1566$ $0.8554$
Query Variants Generation Prompts : Number of Documents Retrieved : Retrieval $0.0001$ $1$ $0.1143$ $0.7363$
Context Words Length Limits : Answer Generation Prompts $0.0001$ $1$ $0.0491$ $0.8253$
Query Variants Generation : Number of Documents Retrieved $0.0000$ $1$ $0.0226$ $0.8808$
Number of Documents Retrieved $0.0000$ $1$ $0.0118$ $0.9137$
Residual $0.0978$ $77$
@@ 3.3. $N$-Way ANOVA Analysis
The ANOVA analysis was performed on the same 96 parameter configurations used in the GoP evaluation. Since the Number of Query Variants and Re-ranker settings were held constant (fixed at eight and Pointwise, respectively), they were excluded from the ANOVA model and are not shown in Table 2. The results presented in Table 2 indicate that Question Augmentation is the most influential individual component affecting system performance. Furthermore, we observed a statistically significant interaction effect between Query Variants Generation Prompts and Retrieval. While neither of these components had a significant effect on their own, their interaction suggests that specific combinations of prompts and retrieval methods can jointly influence performance.
@@ 3.4. Post-ANOVA Improvement: Hypothetical Answer Generation
To further explore improvements in this component, we adopted a hypothetical answer generation technique inspired by HyDE (Gao et al., 2023). This method introduces GRAG, a variant of our RAG system that employs hypothetical answer generation as the Question Augmentation strategy, replacing the “None” setting used in the champion configuration. Due to time constraints, we did not re-run the full GoP or ANOVA tests for this single parameter change. Instead, we conducted a focused comparison between the original champion configuration and GRAG. This ad-hoc evaluation was performed using a sample of 100 questions and measured performance in terms of both Relevance and Faithfulness.
Table 3. Comparison of the champion configuration and GRAG across 100 sampled questions. The table reports how often each system achieved higher or tied Relevance and Faithfulness scores on a per-question basis.
@@
We first compared the two systems at the aggregate level using the average Relevance score. Both the original champion and the GRAG variant achieved the same average score of 1.75. To gain deeper insights, we conducted a single-question analysis. As shown in Table 3, the GRAG configuration outperformed the original on 8 questions and underperformed on 7 in terms of Relevance. For Faithfulness, it scored higher on 12 questions and lower on 14. Given the time constraints and the relatively balanced performance across metrics, we prioritized Relevance as the primary evaluation criterion. Based on this, we selected the GRAG variant as an improvement over the original champion system.
@@
In addition to performance considerations, GRAG also offered improved efficiency. The ANOVA results indicated that Context Length Limitation had no significant impact on evaluation metrics. However, longer contexts substantially increased LLM inference time. To reduce latency during the live event, we limited the input context length to 10k tokens, even though the original champion achieved its best performance with 15k tokens. This decision was further supported by the runner-up configuration, which demonstrated comparable performance with lower computational cost.
@@ 4. Conclusion and Future Work
Through a systematic evaluation of our internal runs – usingDataMorgana to generate synthetic datasets and applying a combination of GoP and $n$-way ANOVA – we efficiently identified the most cost-effective combination of the component parameters to maximize system performance within the time constraints of the LiveRAG 2025 Challenge. This combined approach also enabled us to prioritize the most impactful components and methodically test and refine them, resulting in a well-optimized final configuration for our RAG system. Based on this evaluation, we selected GRAG as our submitted run. The system achieved a Relevance score of 1.199 and a Faithfulness score of 0.477, ranking third on the private LiveRAG 2025 leaderboard and selected as one of the top four finalists based on the final manual evaluation.
@@ 5. Ethical Considerations
It is also worth noting that we have not studied the unintended bias that may get amplified by the use of LLMs in the different stages of GRAG, including the generation of hypothetical answer, query variants, and re-ranking.
@@ Acknowledgements.
This research was conducted by the ARC Centre of Excellence for Automated Decision-Making and Society (ADM+S, Grant #CE200100005), and funded fully by the Australian Government through the Sponsor Australian Research Council https://www.arc.gov.au/ and was undertaken with the assistance of computing resources from RACE (RMIT AWS Cloud Supercomputing). This work was conducted on the unceded lands of the Woi wurrung and Boon wurrung language groups of the eastern Kulin Nation. We pay our respect to Ancestors and Elders, past and present, and extend that respect to all Aboriginal and Torres Strait Islander peoples today and their connections to land, sea, sky, and community.
@@ A.1.2. Medium Prompt
Inspired by Min et al. (2025), although the task in this paper is not totally aligned with query variants generation, we adapted it into a query generation task.
@@ A.3.3. Advanced Prompt
We adapted Tamber et al.’s prompt Tamber et al. (2025), used for their LLM-as-a-judge approach – FaithJudge. [^8] https://github.com/vectara/FaithJudge

### v2

@@ Abstract.
This paper presents the RMIT–ADM+S winning system in the SIGIR 2025 LiveRAG Challenge. Our Generation-Retrieval-Augmented Generation (G-RAG) approach generates a hypothetical answer that is used during the retrieval phase, alongside the original question. G-RAG also incorporates a pointwise large language model (LLM)-based re-ranking step prior to final answer generation. We describe the system architecture and the rationale behind our design choices. In particular, a systematic evaluation using the Grid of Points approach and $N$-way ANOVA enabled a controlled comparison of multiple configurations, including query variant generation, question decomposition, rank fusion strategies, and prompting techniques for answer generation. The submitted system achieved the highest Borda score based on the aggregation of Coverage, Relatedness, and Quality scores from manual evaluations, ranking first in the SIGIR 2025 LiveRAG Challenge.
@@
journalyear: 2025copyright: ccconference: ; July 17, 2025; Padua, Italybooktitle: SIGIR 2025 LiveRAG Challenge, July 17, 2025, Padua, Italydoi: 10.48550/arXiv.2506.14516
@@ 1. Introduction
Evaluation campaigns such as the SIGIR 2025 LiveRAG Challenge (Carmel et al., 2025) provide a structured and standardized setting for researchers and practitioners to develop and evaluate different Retrieval-Augmented Generation (RAG) approaches on a common dataset using shared metrics. These campaigns enable fair comparisons between systems in a controlled environment. LiveRAG 2025, the first edition of the challenge, required participants to develop RAG systems using the Falcon3-10B-Instruct model (Team, 2024) for final answer generation, thereby standardizing the generation component across all submissions. This constraint allows for a more focused evaluation of the retrieval components and prompts, as the generation model is fixed and does not introduce variability in the results. [^1] Other open-weight LLMs, such as the Llama models, were allowed in other components of the system, but only up to their 10B versions.
In this context, the RMIT–ADM+S team submitted G-RAG, a system selected through an internal evaluation process using a Grid of Points (GoP) approach and $N$-way ANOVA (Albahem et al., 2021). [^2] The team includes members from the RMIT Centre for Human-AI Interaction (CHAI) and the ARC Centre of Excellence for Automated Decision-Making and Society (ADM+S). The submitted system ranked first in the LiveRAG 2025 Challenge based on the aggregated manual evaluation scores, and also achieved the highest individual scores in the Coverage (1.61), Relatedness (1.88), and Quality (1.67) metrics. All scores were measured on a 0-2 Likert scale (Carmel et al., 2025). Our submission builds on previous work and integrates several components: (1) hypothetical answer generation prior to retrieval; (2) large language model (LLM)-based query variant generation (Alaofi et al., 2023; Ran et al., 2025); (3) LLM-based re-ranking (Sun et al., 2025); and (4) answerability estimation in RAG systems (Pathiyan Cherumanal et al., 2024). We also developed supporting infrastructure for dynamic cloud-based resource allocation and LLM deployment in the AWS environment, enabling more efficient resource usage.
@@
The source code for both the system and the evaluation framework are publicly available at https://github.com/rmit-ir/G-RAG-LiveRAG.
2. G-RAG System Architecture
Figure 1 illustrates the pipeline of our proposed approach and the system used in our submitted run. In addition to the retrieval and answer generation stages of a RAG system, G-RAG includes two additional stages: question augmentation and re-ranking. For all components involving LLMs, we used the same fixed open-weight model: Falcon3-10B-Instruct. The prompts used throughout the system are provided in Appendix A.
Figure 1. Pipeline of our G-RAG approach. Components used in the final selected run are shown with solid borders; components analyzed during in-house evaluation are shown with dashed borders.
@@ 3.1.1. Dataset Preparation
Table 1. Configuration used to generate the synthetic test collection with DataMorgana.
@@ 3.2. Grid of Points
Table 2. $N$-way ANOVA results. This analysis utilized the average-level LLM-based evaluation method described in Section 3.1.3. A component has a statistically significant impact if its probability value, PR(>F), is less than 0.05. The header lines are as follows: Component: The component being analyzed, refer to Section 3.1.2. SS: Sum of Squares. MS: Mean Square. DF: Degrees of Freedom. F: F-statistic. PR(>F): Probability value greater than F, i.e., $p$-value. $\omega_{p}^{2}$: Partial Omega Squared indicating effect size.
Component SS DF MS F PR(>F) $\omega_{p}^{2}$
Question Augmentation $0.3193$ $2$ $0.1596$ $125.6920$ $<0.0001$* $0.7571$
Query Variants Generation Prompts : Retrieval $0.0062$ $1$ $0.0062$ $4.9099$ $0.0297$* $0.0472$
Query Variants Generation Prompts $0.0030$ $1$ $0.0030$ $2.3892$ $0.1263$ $0.0173$
Context Words Length Limits $0.0025$ $1$ $0.0025$ $2.0052$ $0.1608$ $0.0126$
Retrieval $0.0024$ $1$ $0.0024$ $1.8631$ $0.1762$ $0.0108$
Number of Documents Retrieved : Retrieval $0.0020$ $1$ $0.0020$ $1.6100$ $0.2083$ $0.0077$
Answer Generation Prompts $0.0014$ $1$ $0.0014$ $1.1061$ $0.2962$ $0.0013$
Context Words Length Limits : Question Augmentation $0.0008$ $2$ $0.0004$ $0.3261$ $0.7227$ $-0.0171$
Context Words Length Limits : Answer Generation Prompts : Question Augmentation $0.0004$ $2$ $0.0002$ $0.1566$ $0.8553$ $-0.0215$
Answer Generation Prompts : Question Augmentation $0.0004$ $2$ $0.0002$ $0.1566$ $0.8554$ $-0.0215$
Query Variants Generation Prompts : Number of Documents Retrieved : Retrieval $0.0001$ $1$ $0.0001$ $0.1143$ $0.7363$ $-0.0113$
Context Words Length Limits : Answer Generation Prompts $0.0001$ $1$ $0.0001$ $0.0491$ $0.8253$ $-0.0122$
Query Variants Generation : Number of Documents Retrieved $0.0000$ $1$ $0.0000$ $0.0226$ $0.8808$ $-0.0125$
Number of Documents Retrieved $0.0000$ $1$ $0.0000$ $0.0118$ $0.9137$ $-0.0127$
Residual $0.0978$ $77$ $0.0013$
@@ 3.3. $N$-Way ANOVA Analysis
The ANOVA analysis was performed on the same 96 parameter configurations used in the GoP evaluation. Since the Number of Query Variants and Re-ranker settings were held constant (fixed at eight and Pointwise, respectively), they were excluded from the ANOVA model and are not shown in Table 2. The results presented in Table 2 indicate that Question Augmentation is the most influential component affecting system performance. The dominant influence of Question Augmentation is evident not only in its statistical significance ($p$-value$<0.0001$), but also in its effect size. The Partial Omega Squared ($\omega_{p}^{2}=0.7571$) underscores its practical importance, accounting for the majority of explained variance. In contrast, the negative $\omega_{p}^{2}$ for some components, which occur when an F-statistic is less than $1.0$, signify the absence of effect (Kroes and Finley, 2023). Furthermore, we observed a statistically significant interaction effect between Query Variants Generation Prompts and Retrieval. While neither of these components had a significant effect on their own, their interaction suggests that specific combinations of prompts and retrieval methods can jointly influence performance.
@@ 3.4. Post-ANOVA Improvement: Hypothetical Answer Generation
To further explore improvements in this component, we adopted a hypothetical answer generation technique inspired by HyDE (Gao et al., 2023). This method introduces G-RAG, a variant of our RAG system that employs hypothetical answer generation as the Question Augmentation strategy, replacing the “None” setting used in the champion configuration. Due to time constraints, we did not re-run the full GoP or ANOVA tests for this single parameter change. Instead, we conducted a focused comparison between the original champion configuration and G-RAG. This ad-hoc evaluation was performed using a sample of 100 questions and measured performance in terms of both Relevance and Faithfulness.
Table 3. Comparison of the champion configuration and G-RAG across 100 sampled questions. The table reports how often each system achieved higher or tied Relevance and Faithfulness scores on a per-question basis.
@@
We first compared the two systems at the aggregate level using the average Relevance score. Both the original champion and the G-RAG variant achieved the same average score of 1.75. To gain deeper insights, we conducted a single-question analysis. As shown in Table 3, the G-RAG configuration outperformed the original on 8 questions and underperformed on 7 in terms of Relevance. For Faithfulness, it scored higher on 12 questions and lower on 14. Given the time constraints and the relatively balanced performance across metrics, we prioritized Relevance as the primary evaluation criterion. Based on this, we selected the G-RAG variant as an improvement over the original champion system.
@@
In addition to performance considerations, G-RAG also offered improved efficiency. The ANOVA results indicated that Context Length Limitation had no significant impact on evaluation metrics. However, longer contexts substantially increased LLM inference time. To reduce latency during the live event, we limited the input context length to 10k tokens, even though the original champion achieved its best performance with 15k tokens. This decision was further supported by the runner-up configuration, which demonstrated comparable performance with lower computational cost.
@@ 4. Conclusion and Future Work
Through a systematic evaluation of our internal runs – usingDataMorgana to generate synthetic datasets and applying a combination of GoP and $n$-way ANOVA – we efficiently identified the most cost-effective combination of the component parameters to maximize system performance within the time constraints of the LiveRAG 2025 Challenge. This combined approach also enabled us to prioritize the most impactful components and methodically test and refine them, resulting in a well-optimized final configuration for our RAG system. Based on this evaluation, we selected G-RAG as our submitted run. The system achieved a Relevance score of 1.199 and a Faithfulness score of 0.477, ranking third on the private LiveRAG 2025 leaderboard and selected as one of the top four finalists based on the final manual evaluation (Carmel et al., 2025). G-RAG was ultimately announced as the winning system, validating the effectiveness of our design and evaluation strategy.
@@ 5. Ethical Considerations
It is also worth noting that we have not studied the unintended bias that may get amplified by the use of LLMs in the different stages of G-RAG, including the generation of hypothetical answer, query variants, and re-ranking.
@@ Acknowledgements.
We thank the SIGIR 2025 LiveRAG Challenge organizers for the opportunity to participate and their support, and the reviewers for their helpful feedback. This research was conducted by the ARC Centre of Excellence for Automated Decision-Making and Society (ADM+S, Grant #CE200100005), and funded fully by the Australian Government through the Sponsor Australian Research Council https://www.arc.gov.au/ and was undertaken with the assistance of computing resources from RACE (RMIT AWS Cloud Supercomputing). This work was conducted on the unceded lands of the Woi wurrung and Boon wurrung language groups of the eastern Kulin Nation. We pay our respect to Ancestors and Elders, past and present, and extend that respect to all Aboriginal and Torres Strait Islander peoples today and their connections to land, sea, sky, and community.
@@ A.1.2. Medium Prompt
Inspired by Min et al. (2025), although the task in this paper is not completely aligned with query variants generation, we adapted it into a query generation task.
@@ A.3.3. Advanced Prompt
We adapted Tamber et al.’s prompt (Tamber et al., 2025), used for their LLM-as-a-judge approach – FaithJudge. [^8] https://github.com/vectara/FaithJudge